OpenStudy (anonymous):

1 / 1 - 3^2x = 4 find the x

OpenStudy (anonymous):

do you mind, to give the steps ?

OpenStudy (anonymous):

did u mean 1/[1-3^(2x)]=4 or (1/1) -3^(2x)=4 or (1/1)-(3^2)x =4